OpenStudy (anonymous):

MEDAL AND FAN!! A hyperbola with a horizontal transverse axis contains the point at (4, 3). The equations of the asymptotes are y - x = 1 and y + x = 1. Write an equation for the hyperbola.

OpenStudy (anonymous):

@ganeshie8 please help

OpenStudy (anonymous):

@math-geek @hartnn help

OpenStudy (anonymous):

@IMStuck can u help me

OpenStudy (anonymous):

@AravindG

OpenStudy (imstuck):

Give me a sec to work on this one...hyperbolas are stinkers! ; )

OpenStudy (anonymous):

sure thanks in advance

OpenStudy (anonymous):

umm...u there???

OpenStudy (imstuck):

I am and I am really struggling to find the a and the b here.

OpenStudy (imstuck):

The best I can do is this:\[\frac{ x ^{2} }{ 1 }-\frac{ (y-1)^{2} }{ 1 }=1\]

OpenStudy (anonymous):

okay thnx

OpenStudy (imstuck):

ACtually I think I DID get it...\[\frac{ x ^{2} }{ 4 }-\frac{ (y-1)^{2} }{ 3 }=1\]

OpenStudy (imstuck):

Can you check it?

OpenStudy (anonymous):

i checked it it is incorrect but i got it now so nvr mind

OpenStudy (imstuck):

what is it!? Please tell me so I can find out how to find a and b. Please!

OpenStudy (anonymous):

its very similar to my ques

OpenStudy (imstuck):

OMG thank you for that website! I love these informational sites; they are such a great help!

OpenStudy (anonymous):

ur welcum

OpenStudy (mathmate):

We're given y-x=1 and y+x=1, which are lines that intersect at (0,1) at 90 degrees. This means that the "rectangle" is a square, or a=b. c (distance of the centre to the foci) is therefore \(\sqrt2 a\). The general equation of a hyperbola with centre (h,k) is given by: \(\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\) since the hyperbola has a horizontal axis. The conic has a centre of (0,1), so the equation reduces to: \(\frac{x^2}{a^2}+\frac{(y-1)^2}{b^2}=1\) In addition, from the angles of the two asymptotes, we further deduced that a=b. The equation is therefore: \(\frac{x^2}{a^2}+\frac{(y-1)^2}{a^2}=1\) Now we'll substitute (4,3) into the equation to solve for a. \(\frac{4^2}{a^2}+\frac{(3-1)^2}{a^2}=1\) => \(a^2=16-4=12\) or \(a^2=12\) The final equation is therefore: \(\frac{x^2}{12}+\frac{(y-1)^2}{12}=1\)